OpenStudy (anonymous):

A. 60 B. 57 C. 29 D. 33

OpenStudy (blurbendy):

What is the question?

OpenStudy (johnweldon1993):

And I take it we're to find the measure of the angle? You have an angle...a side adjacent to that angle...and the hypotenuse....use the COS function \[\cos(x) = \frac{ 11 }{ 20 }\] Now you want to find the angle...so use the ARCCOS function \[x = \cos^{-1} \frac{ 11 }{ 20 }\] *make sure your calculator is in degree mode*

OpenStudy (anonymous):

I dont have a calculator with cos

OpenStudy (anonymous):

You can calculate it in Google ! Type : arccos (11/20)

OpenStudy (anonymous):

Thank you. I will try!

OpenStudy (anonymous):

arccos(11 / 20) = 0.988432089 rad ?????

OpenStudy (anonymous):

Yes !

OpenStudy (anonymous):

it is not one of my choices

OpenStudy (anonymous):

These are multiple choice questions

OpenStudy (johnweldon1993):

Type in "arccos (11/20) in degrees"

OpenStudy (johnweldon1993):

Did you get it @Joshyface ?

OpenStudy (anonymous):

57

OpenStudy (johnweldon1993):

That would be correct! Good work!